OpenStudy (insa):

A basketball has a surface area of 1810cm^2. Find the radius and the volume of air in it

OpenStudy (whpalmer4):

Okay, you'll need the formula for the surface area of a sphere: \[A = 4 \pi r^2\] You know the surface area, so you should be able to solve that equation for \(r\) in terms of \(A\), plug in the value of \(A\) and find \(r\). Then plug your value for \(r\) into the equation for the volume of a sphere:\[V = \frac{4}{3}\pi r^3\]to find the volume of the air inside the basketball (we are apparently assuming that the basketball has a shell which occupies no space!)

OpenStudy (whpalmer4):

Don't forget to determine the proper unit for your answer!